OpenStudy (anonymous):

Can someone help me on these two questions 3/2x-15=-12 and 4x+1/2(12x+2)=3

OpenStudy (allank):

for the first, add 15 to both sides 3/2x=-12+15 3/2x=3 x=3*2/3=2

OpenStudy (anonymous):

For the second one: multiply that 1/2 through the parentheses so you get 4x+(6x+1)=3 or 4x+6x+1=3 Add the like terms: 10x+1=3 Subtracted 1 from both sides: 10x=2 Divide both sides by 10: x=2/10 Reduce: x=1/5

OpenStudy (allank):

@mdttheory nice teamwork

OpenStudy (anonymous):

I'm confused on the first one.

OpenStudy (allank):

have to go. do explain, @mdttheory

OpenStudy (anonymous):

3/2x-15=-12 Add 15 to both sides to get you: 3/2x=3 Now what you basically have is (3)(1/2)(x)=3 which is the same thing as 3/2x=3 You can get rid of the 3 by dividing both sides by 3 You can get rid of the 1/2 by multiplying both sides by 2

OpenStudy (anonymous):

@mdttheory- would the answer for the the second one be x= 2/10 and x=1/5?

OpenStudy (anonymous):

2/10=1/5 So, yeah, both are 'correct' as they are the same thing but 1/5 is the simplified version of the answer which most teachers ask you to find.

OpenStudy (anonymous):

Ohh ok thank you! =)

OpenStudy (anonymous):

For the second one i added 15 on both sides and i didn't get exactly what to do to get rid of the fractions

OpenStudy (anonymous):

Would the answer be x=2

OpenStudy (anonymous):

Well, you can either break the fraction up into its parts like I did or you can treat it like a whole number in itself. What I mean to say is if it were 5x=3 you would simply divide both sides by 5, right? Do the same thing with the fraction. If you have 3/2x=3 you just divide both sides by 3/2. Note: when you divide by a fraction it's the same thing as multiplying by its inverse or: x=(3)(2/3) <- Just flipped over the fraction. So, yes, the 3's cancel and you get 2.

OpenStudy (anonymous):

Ohhhh ok, I get it now. Thank you SO much mdttheory!! Your a big help =).

OpenStudy (anonymous):

np, good luck

OpenStudy (anonymous):

Thank you =)